# ChartMoE: Mixture of Expert Connector for Advanced Chart Understanding

### v1

@@ 4.1 Implementation Details
The ChartMoE builds on InternlmXC-v2 Dong et al. (2024), integrating ViT-Large Radford et al. (2021) as the visual encoder and InternLM-v2 Cai et al. (2024) as the language model. We replace the MLP with MoE architecture as the connector. During the alignment stage, we train the connector parameters and keep the visual encoder and LLM parameters fixed for 1 epoch. The learning rate is set to 5e-5 with a warmup phase covering the first 1% of training steps. In the supervised fine-tuning stage, we continue training the connector while employing LoRA to update the LLM parameters with the rank of 64. The learning rate is adjusted to 1e-5 for the high-quality knowledge learning period and 5e-5 for the chart-specific annealing tuning period. The weight decay is 0.1 for all stages. We use the cosine annealing learning rate schedule. The global batch size is set to 64 for all stages. The training process is conducted on A100-40G GPU, with the alignment stage taking approximately 240 GPU hours, the knowledge learning stage taking 138 GPU hours, and the chart-specific annealing tuning stage taking around 76 GPU hours.
4.2 Evaluation Metrics
ChartQA Masry et al. (2022) test split consists of 1,250 questions in both human and augmented parts. The charts are three common chart types and are sourced from the real world. It features a variety of human-crafted questions and answers to evaluate models’ understanding, reasoning, and data extraction skills. ChartQA adopts relaxed accuracy, which is highlighted shortcomings by recent studies Chen et al. (2024); Xu et al. (2023), such as simplistic string matching and direct float conversion. Therefore, we improve it by 1) using regular expression matching to extract number values, 2) optimizing string matching for short answers, and 3) demonstrating model performance under various relaxed margins. We adopt it for all experiment results in this paper.
ChartFC Akhtar et al. (2023a) & ChartCheck Akhtar et al. (2023b) They use accuracy to verify whether the claim aligns with the input chart, marking a significant advancement in chart recognition and reasoning abilities. This also identifies the potential hallucinations in chart-related contexts. The ChartFC test set has 1,591 questions, and the ChartCheck test set has two splits, containing 937 questions and 981 questions.
4.3 Comparative Models
Table 2: The relaxed accuracy (%) performance on ChartQA. Ada.: Adaptive input resolution. *: Multi-scale image feature, 448$\times$9 in DocOwl. $\dagger$: Employing token merging to reduce computational overhead.
@@ 4.4 Main Results
ChartQA. Tab.2 presents detailed comparisons of ChartMoE on ChartQA. ChartMoE significantly improves the baseline (InternlmXC-v2) performance (72.00% vs. 84.64%, +12.64$\uparrow$ in Acc.@0.05). Compared to previous SOTA (TinyChart+PoT @768 pixel), ChartMoE consistently surpasses it across all metrics. The PoT effectively enhances the model’s mathematical reasoning capabilities, which is a common shortfall in current MLLMs. ChartMoE integrates better with PoT, indicating that it accurately extracts fundamental elements from charts. The questions in Human part tend to be more computationally complex and challenging, and ChartMoE shows more significant improvement in this area, especially after incorporating PoT. Notably, our error analysis in the Augmented part reveals that many errors stem from limitations of the evaluation criteria, i.e., string matching. For instance, it is marked incorrect if the prediction is It is between 2003 and 2005 and the ground truth is (2003, 2005). Forcing performance improvement on it may lead to model overfitting.
ChartBench. Tab. 3 presents detailed comparisons of ChartMoE on ChartBench. None of the models, including our ChartMoE, undergo supervised finetuning on the ChartBench trainset to ensure fair experimental comparison. Chart-specific models typically underperform due to their limited generalization ability, which prevents them from effectively managing charts with sparse data point annotations ($<10$%). Deplot shows a distinct advantage over these types of models (27.62%) with the assistance of LLaVA-v1.6. The baseline (InternlmXC-v2) demonstrates strong generalization on ChartBench (48.41%), which may benefit from pre-training instructions designed for unannotated charts. Without additional design, ChartMoE improves the baseline performance comprehensively (48.41% vs. 51.67%), especially on extra chart types (39.72% vs. 55.58%, +15.86$\uparrow$).
@@ 5.1 Model Architecture Ablation
Effect of Number of Experts and Activated Experts. As shown in rows 3-4 of Tab. 6, we compare ChartMoE configurations with 4 and 8 experts, keeping 2 experts activated. The 8 experts are initialized in pairs using the 4 methods illustrated in Fig. 2c. ChartMoE achieves 78.76% in Acc.@0.05 with 4 experts, which is slightly higher than the 78.60% achieved with 8 experts, showing a marginal increase of +0.16%. In rows 4-5, we compare the performance of configurations with 2 and 4 activated experts, finding similar results: 78.60% vs. 78.64% in Acc.@0.05. This analysis indicates that simply increasing the number of experts or activated experts does not necessarily lead to better performance. The configuration with 4 experts and 2 activated experts effectively balances complexity and performance, making it a suitable choice for ChartMoE.
@@ 5.2 Training Strategy Ablation
4 2 Random 59.68 86.72 73.20
4 2 Aligned Co-Upcycle 64.96 90.00 77.48
4 2 Ours 67.92 89.60 78.76
8 2 Ours 67.20 90.00 78.60
8 4 Ours 67.68 89.60 78.64
Table 6: Ablation study on the training strategy, about the alignment stage, high-quality knowledge learning period and chart-specific anneal tuning period.
@@
Table 8: Ablation study on alignment pertaining tasks. We adopt different alignment tasks for baseline (linear connector) and further conduct supervised fine-tuning on the ChartQA train set.
@@
Effect of Alignment Strategy. As shown in rows 1-3 of Tab. 6, translating the chart image into structural text formats such as table, JSON, and code during the alignment stage significantly enhances performance in downstream chart understanding tasks. After applying table-JSON-code alignment, the model achieves 77.20% in Acc.@0.05, representing a notable improvement of +1.52$\uparrow$. When combined with our proposed MoE architecture, performance further increases to 78.76%, a total gain of +3.08$\uparrow$ in Acc.@0.05.
@@ 5.3 In-depth Analysis
Effect of Alignment Task. As shown in Tab.8, we explore various alignment tasks based on the baseline (linear connector). After alignment, the performance significantly declines compared to the vanilla version. However, the aligned model exhibits a substantial improvement after supervised finetuning on the ChartQA train split, which is consistent with previous observations Meng et al. (2024); Yan et al. (2024). Specifically, alignment based on JSON and code exhibits remarkable improvement over the one based on the table.
@@
5.4 Expert Distribution Visualiztion
As shown in Fig. 5 and Fig. 6, we visualize the expert distribution in the MoE connector on the ChartBench test set. We designate the vanilla connector as E0, while E1-3 corresponds to connectors aligned with tables, JSON, and code. As depicted in Fig. 5, the trend is consistent across different chart types, with E0 and E3 being the most frequently selected connectors. The expert selection shows no extreme bias, as even the least chosen, E1, accounts for over 10%.
We further visualize the expert selection for each image token, revealing the preferences of each expert. As shown in Fig. 6, E0 is the primary choice for background tokens, explaining its dominance in Fig. 5. E1 and E2 are more frequently chosen by tokens from titles, axis labels, or legends, as these elements are commonly found in tables and JSON files. ChartMoE tends to use E3 to focus on the data points and visual elements within the chart, e.g., data points on the line, digital text, and edges in a node chart. These components are essential for accurately re-drawing the charts.
@@ 6 Conclusion
We introduce ChartMoE, a multi-task aligned and instruction-tuned MLLM designed for complex chart understanding and reasoning. We replace the linear connector with the MoE architecture and initialize each expert with parameters derived from different alignment tasks. We further present the ChartMoE-Align dataset, a synthetic collection of over 900K table-json-code-chart quadruples, to facilitate alignment training across different experts. This approach preserves the strengths of each alignment task, ensuring efficient training and superior model performance. ChartMoE outperforms the previous state-of-the-art on several benchmarks by a large margin and excels in real-world applications such as chart question answering, translation, and editing.
@@ A.1 Top-2 experts distribution
(c) ChartMoE
Figure 7: The distribution of Top-2 experts after supervised fine-tuning with three expert initialization methods. We calculate the proportion of the top 2 experts selected by the MoE connector on the Test split of ChartBench.
@@ A.2 Summary of hyperparameter settings
The training process of our ChartMoEis structured into three distinct phases: Alignment Pretraining, High-Quality Knowledge Learning and Chart-Specific Annealing Tuning. Table 9 provides a comprehensive overview of the hyperparameter configurations employed during each of these training stages.
@@
Configuration Alignment Pretraining High-Quality Knowledge Learning Chart Specific Annealing Tuning
@@
Peak Learning Rate $5e^{-5}$ $1e^{-5}$ $5e^{-5}$
@@
Global Batch Size 64 32 32
Gradient Acc. 8 8 8
@@
GPU Hours (A100-40G) 240 138 76
@@ Appendix B Additional Vsualizations of top-1 expert selection
Appendix C Detailed Examples in ChartMoE-Align
@@
Appendix D Additional Capabilities on Chart Editing and Highlighting
Leveraging fine-grained alignment, particularly in translating input charts into corresponding Python code, ChartMoE demonstrates advanced capabilities in editing and highlighting key elements. As illustrated in Fig. 11, ChartMoE effectively modifies input charts based on user instructions, such as converting a vertical bar chart into a horizontal one, transforming a pie chart into a bar chart, and adjusting the values of specific elements.
ChartMoE also excels at adding auxiliary lines and highlighting elements within charts, facilitating the rapid extraction of valuable insights. For instance, in Fig. 12, ChartMoE draws a horizontal line through the lowest data point and both horizontal and vertical lines through each data point, aiding in the precise determination of values. Additionally, in Fig. 13, ChartMoE highlights the maximum data point using two different methods: employing a larger green scatter with a prominent marker and applying a bright color to make it stand out.

### v3

@@ 4.1 Implementation Details
During the alignment stage, we train the connector parameters and keep the visual encoder and LLM parameters frozen. In the supervised fine-tuning stage, we continue training the MoE connector while employing LoRA to update the LLM parameters. All training processes are conducted on 4 $\times$ A100-40G GPUs. You can refer to Appendix A.2 for more details.
Table 2: The relaxed accuracy (%) performance on ChartQA. Ada.: Adaptive input resolution. *: Multi-scale image feature, 448$\times$448 in DocOwl. $\dagger$: Employing token merging to reduce computational overhead.
@@
4.2 Evaluation Metrics
ChartQA Masry et al. (2022) test split consists of 1,250 questions in both human and augmented parts. The charts are three common chart types and are sourced from the real world. It features a variety of human-crafted questions and answers to evaluate models’ understanding, reasoning, and data extraction skills. ChartQA adopts relaxed accuracy, which is highlighted shortcomings by recent studies Chen et al. (2024); Xu et al. (2023), such as simplistic string matching and direct float conversion. Therefore, we improve it by 1) using regular expression matching to extract number values, 2) optimizing string matching for short answers, and 3) demonstrating model performance under various relaxed margins. We adopt it for all experiment results.
ChartFC Akhtar et al. (2023a) & ChartCheck Akhtar et al. (2023b) adopt accuracy to verify whether the claim aligns with the input chart, marking a significant advancement in chart recognition and reasoning abilities. This identifies the potential hallucinations in chart-related contexts. The ChartFC test set has 1,591 questions, and the ChartCheck test set has two splits, containing 937 questions and 981 questions.
4.3 Comparative Models
@@ 4.4 Main Results
ChartQA. Tab.2 presents detailed comparisons of ChartMoE on ChartQA. ChartMoE significantly improves the baseline (InternlmXC-v2) performance (72.00% vs. 84.64%, +12.64$\uparrow$ in Acc.@0.05). Compared to previous SOTA (TinyChart+PoT @768 pixel), ChartMoE consistently surpasses it across all metrics. The PoT effectively enhances the mathematical reasoning capabilities, which is a common shortfall in current MLLMs. ChartMoE integrates better with PoT, indicating that it accurately extracts fundamental elements from charts. ChartMoE shows more significant improvement on Human part, especially after incorporating PoT, where the questions are more computationally complex and challenging. Notably, our error analysis in the Augmented part reveals that many errors stem from limitations of the evaluation criteria, i.e., string matching. For instance, it is marked incorrect if the prediction is It is between 2003 and 2005 and the ground truth is (2003, 2005). Forcing performance improvement may lead to model overfitting.
ChartBench. Tab. 3 presents detailed comparisons of ChartMoE on ChartBench. None of the models, including our ChartMoE, undergo supervised fine-tuning on the ChartBench trainset to ensure fair experimental comparison. Chart-specific models typically underperform due to limited generalization, which fails to manage the annotated charts effectively ($<10$%). Deplot shows a distinct advantage over these types of models (27.62%) with the assistance of LLaVA-v1.6. The baseline (InternlmXC-v2) demonstrates strong generalization on ChartBench (48.41%), which may benefit from pre-training instructions designed for unannotated charts. Without additional design, ChartMoE improves the baseline performance comprehensively (48.41% vs. 51.67%), especially on extra chart types (39.72% vs. 55.58%, +15.86$\uparrow$).
@@ 5.1 Model Architecture Ablation
Effect of Number of Experts and Activated Experts. As shown in rows 3-4 of Tab. 6, we compare ChartMoE configurations with 4 and 8 experts, keeping 2 experts activated. The 8 experts are initialized in pairs using the 4 methods illustrated in Fig. 2c. ChartMoE achieves 78.76% in Acc.@0.05 with 4 experts, which is slightly higher than the 78.60% achieved with 8 experts, showing a marginal increase of +0.16%. In rows 4-5, we compare the performance of configurations with 2 and 4 activated experts, finding similar results: 78.60% vs. 78.64% in Acc.@0.05. This analysis suggests that merely increasing the number of experts or the activation of experts does not guarantee improved performance. The configuration with 4 experts and 2 activated experts effectively balances complexity and performance, making it a suitable choice for ChartMoE.
@@ 5.2 Training Strategy Ablation
4 2 Random Init. 59.68 86.72 73.20
4 2 Random Align 62.32 88.88 75.60
4 2 Co-Upcycle Init. 64.96 90.00 77.48
4 2 Diversely Align 67.92 89.60 78.76
8 2 Diversely Align 67.20 90.00 78.60
8 4 Diversely Align 67.68 89.60 78.64
Table 6: Ablation study on the proposed training strategy and connector architecture on the alignment, high-quality knowledge learning, and chart-specific anneal tuning stages.
@@
Table 8: Ablation study on alignment pre-training tasks. We adopt different alignment tasks for baseline (linear connector) and further conduct supervised fine-tuning on the ChartQA train set.
@@
Effect of Alignment Strategy. As shown in rows 1-3 of Tab. 6, translating the chart image into structural text formats such as table, JSON, and code during the alignment stage significantly enhances performance in downstream chart understanding tasks. After applying table-JSON-code alignment, the model achieves 77.20% in Acc.@0.05, representing a notable improvement (+1.52$\uparrow$). When combined with our proposed MoE connector, performance further increases to 78.76%, a total gain of +3.08$\uparrow$ in Acc.@0.05.
@@ 5.3 In-depth Analysis
Effect of Alignment Task. As shown in Tab. 8, we explore various alignment tasks based on the linear connector. After alignment, the performance on ChartQA declines compared to the baseline. However, the aligned model exhibits a substantial improvement after supervised fine-tuning on the ChartQA train split, which is consistent with previous observations Meng et al. (2024); Yan et al. (2024). Specifically, the JSON and code tasks exhibit remarkable improvement over the table.
@@
Figure 7: The performance on the general VQA tasks (MME Fu et al. (2023)). With supervised fine-tuning on extensive chart-structured data, the directly tuned IXC-v2 shows a significant performance drop, while ChartMoE maintains a satisfying performance by keeping the vanilla connector as the expert in MoE.
Figure 8: The performance on the general VQA tasks (MMBench Liu et al. (2023e)). Please refer to its paper for each task’s details. The observations and conclusions are consistent with the MME benchmark.
Figure 9: The performance with/without bz-loss on ChartQA. Left: The bz-loss leads to more even expert selections. Right: A more balanced distribution does not yield better performance.
Expert Distribution Visualization. As shown in Fig. 56, we visualize the expert distribution in the MoE connector on the ChartBench test set. We designate the vanilla connector as E0, while E1-3 corresponds to connectors aligned with tables, JSON, and code. As depicted in Fig. 5, the trend is consistent across different chart types, with E0 and E3 being the most frequently selected connectors. The expert selection shows no extreme bias, as even the least chosen, E1, accounts for over 10%. We further visualize the expert selection for each image token, revealing the preferences of each expert. As shown in Fig. 6, E0 is the primary choice for background tokens, explaining its dominance in Fig. 5. E1 and E2 are more frequently chosen by tokens from titles, axis labels, or legends, as these elements are commonly found in tables and JSON files. ChartMoE tends to use E3 to focus on the data points and visual elements within the chart, e.g., data points on the line, digital text, and edges in a node chart. These components are essential for accurately re-drawing the charts.
Performance on General Tasks. While ChartMoE is designed to enhance chart understanding, it does not compromise other capabilities, e.g., instruction following and object recognition. In Fig. 78, we show the comparisons of directly fine-tuned InternlmXC-v2 (short for Direct SFT) with data from Tab. 1 and the baseline (short for IXC-v2) on general benchmarks Fu et al. (2023); Liu et al. (2023e). The direct SFT model shows diminished general capabilities. In contrast, ChartMoE preserves it nearly intact by retaining the original connector as one of its experts.
Effect of Balance Loss in MoE. The standard MoE Zoph et al. (2022) employs balanced loss and router z-loss (short for bz-loss) to prevent certain experts from dominating the model training. In Fig. 9, we compare the effects of with and without bz-loss. While bz-loss promotes a more equitable selection of experts, it fails to enhance ChartMoE’s performance further. As shown in Fig. 6, the expert initialization in ChartMoE results in each expert having its own preference for visual token selection (refer to Appendix B for detail). Consequently, the bz-loss might hinder the model’s convergence to the optimal point because the distribution of visual tokens is inherently imbalanced.
@@ 6 Conclusion
We introduce ChartMoE, a multi-task aligned and instruction-tuned MLLM designed for complex chart understanding and reasoning. We replace the linear connector with the MoE architecture and initialize each expert with parameters derived from different alignment tasks. We further present the ChartMoE-Align dataset, a synthetic collection of nearly 1 million table-json-code-chart quadruples, to facilitate alignment training across different experts. This approach preserves the strengths of each alignment task, ensuring efficient training and superior model performance. ChartMoE outperforms the previous state-of-the-art on several benchmarks by a large margin and excels in real-world applications such as chart question answering, translation, and editing. Please refer to Appendix A.3 for the reproducibility statement.
7 ACKNOWLEDGMENTS
This work was supported by the National Key R&D Program of China (2022YFB4701400/470140
2), SSTIC Grant(KJZD20230923115106012, KJZD20230923114916032, GJHZ202402181136040
08), and Beijing Key Lab of Networked Multimedia.
@@ A.1 Top-2 experts distribution
(c) ChartMoE with bz-loss
(d) ChartMoE
Figure 10: The distribution of Top-2 experts after supervised fine-tuning with three expert initialization methods. We calculate the proportion of the top 2 experts selected by the router on the ChartBench.
@@ A.2 Summary of hyperparameter settings
The training process of our ChartMoE is structured into three distinct phases: Alignment Pre-training, High-Quality Knowledge Learning, and Chart-Specific Annealing Tuning. Table 9 provides a comprehensive overview of the hyperparameter configurations employed during each training stage. It should be noted that using flash-attention Dao (2024) can effectively reduce GPU hours.
@@
Configuration Alignment Pre-training High-Quality Knowledge Learning Chart Specific Annealing Tuning
@@
Peak Learning Rate $5e^{-5}$ $5e^{-5}$ $1e^{-5}$
@@
Global Batch Size 256 64 64
Gradient Acc. 16 8 8
@@
GPU Hours (A100-40G) 210 100 56
A.3 Reproducibility Statement
We have included the architecture of ChartMoE in Section 3.1 and the complete training procedure in Section 3.2 and Section 3.3. The training data recipe is listed in Tab. 1 in detail. Hyper-parameter settings are shown in Appendix A.2. We also introduce the generation pipeline for ChartMoE-Align in Section 3.2, and some detailed examples in Appendix C. Furthermore, our ChartMoE-Align dataset and checkpoints of ChartMoE will be released soon on GitHub and Huggingface.
@@ Appendix B Additional Vsualizations of Top-1 Expert Selection
In this section, we randomly sampled images from natural image datasets (LLaVA-CC3M Liu et al. (2023d)) and chart datasets (ChartQA Masry et al. (2022), ChartBench Xu et al. (2023)) to illustrate ChartMoE’s token selection preferences. As shown in Fig. 11, the vanilla expert focuses more on the background, the table expert concentrates on details such as the boundary between the background and the subject, the JSON expert focuses on textures (e.g., maps and objects), and the code expert specializes in curves and trends (e.g., logos and text). Fig. 12 further demonstrates that while the vanilla expert continues to attend to background tokens, critical visual elements are handled by the aligned experts, with the code expert being notably more prominent.
@@
Appendix C Details of ChartMoE-Align
C.1 Overview
ChartMoE-Align is a dataset we introduced for different experts aligning pretraining. It consists of nearly 1 million Chart Table JSON Code quadruples and supports three alignment tasks: Chart to Table, Chart to JSON, and Chart to Code. Unlike other chart datasets, ChartMoE-Align focuses solely on these three fundamental alignment tasks without considering the diversity of instruction tasks.
C.2 Table Data Collection
We primarily collect table data from three sources: the ChartQA training set Masry et al. (2022), the PlotQA training set Methani et al. (2020), and ChartY provided by OneChart Chen et al. (2024).
ChartQA includes 18.3K real-world charts and provides accompanying meta tables. While the charts are of high quality and manually curated, they lack fine-grained attribute annotations and executable plotting code. As a result, we only retained the tables from ChartQA in CSV format.
PlotQA comprises 157K charts, primarily focusing on three common types: line, bar, and pie charts. These charts are generated using Python code with limited formatting and style diversity. Consequently, we did not utilize the charts from PlotQA but retained its 157K tables. These tables originate from sources like World Bank Open Data, Open Government Data, and the Global Terrorism Database, covering statistics on various indicators such as fertility rates, rainfall, coal production, and more across years, countries, and districts.
ChartY is a chart dataset containing 2.7M charts in both Chinese and English proposed by OneChart. Notably, ChartY also includes charts from ChartQA and PlotQA, which we filtered out in ChartMoE-Align. Additionally, ChartY primarily consists of common chart types such as line, bar, and pie charts (or their combinations) and suffers from significant data imbalance. To address this, we sampled a subset to ensure a roughly equal number of charts for each type. As the tables in ChartY are mainly generated by GPT-3.5 based on templates, we ultimately retained 763K samples from this source.
C.3 Pair Data Construction
JSON provides a structured format distinct from CSV, designed to retain chart attributes beyond numerical data, such as title position, font size, element colors, legend styles, and more. We adopt the template provided by ChartReformer Yan et al. (2024) and further enhance it. We add chart type-agnostic attributes like title position and gridlines. For chart type-specific attributes, we aim to remain consistent with ChartReformer’s definitions while accommodating all chart types present in ChartMoE-Align. With this framework, we generate corresponding JSON files for all tables. To extract chart type-specific attributes, we fine-tune a Deplot Liu et al. (2023a) model, leveraging the original chart to extract their properties. Missing attributes are filled in using random sampling to ensure completeness.
Code refers to Python scripts based on matplotlib for rendering the charts. Leveraging the rich attributes defined in the JSON, the code is designed to faithfully represent every attribute to ensure diversity in the resulting charts. During generation, we explicitly specify all default parameters, such as the hexadecimal color codes for each line/bar, default font sizes, text positions, etc. We provide basic code templates for type-agnostic attributes. For type-specific attributes, rules are used to automatically generate the corresponding code.
Chart is produced by executing the generated code. Given the number of table, JSON, and code pairs, we filter out any quadruples with execution errors or warnings during the chart generation process, retaining only valid and error-free samples.
Instruction. Considering the alignment task, we directly employ several templated questions to define the Chart-to-X tasks (X is the ground truth). Ultimately, each quadruple corresponds to three QA pairs. Note that ChartMoE-Align only serves for alignment training to initialize different expert projectors, thus emphasizing the diversity of charts and aligned modalities. To improve model performance and instruction-following, we still require more diverse instructions for supervised fine-tuning to update the MoE connector and LLM.
C.4 Quality Control
We first remove all duplicate entries from the meta table and then eliminate quadruples that cause errors or warnings during rendering. To further assess the quality of ChartMoE-Align, we randomly sample 200 quadruples and ask GPT-4o and annotation experts (with at least three experts reviewing each quadruple) to evaluate the clarity and readability of the charts, as well as the alignment between the charts and table/JSON/code, scoring them as 0 or 1. The results show that nearly all charts are clear, unambiguous, and free from obstructions (GPT-4o: 96.5%, Experts: 99%). Over 90% of the pairs are matching and suitable for instruction tuning (GPT-4o: 91%, Experts: 94.5%).
C.5 Example Visualization
@@
Appendix D Further Discussion
D.1 Contribution of ChartMoE
Some prior work, such as MoE-LLaVA Lin et al. (2024), DeepSeek-VL Lu et al. (2024), and CuMo Li et al. (2024), has employed MoE architectures in MLLMs. However, these approaches all apply MoE to LLMs or ViTs to increase model capacity, introducing a large number of learnable parameters to boost performance. In contrast, our ChartMoE introduces several distinctive innovations:
1) Motivation: Our goal is not to expand model capacity but to enhance the model’s chart comprehension through alignment tasks while preserving performance on other general tasks. Hence, we retain the original connector parameters as one expert initialization manner.
2) Initialization: Unlike previous methods that rely on random or co-upcycle initialization, we leverage multiple alignment tasks for expert (connector) initialization. This approach enables ChartMoE to exhibit remarkable interpretability (Fig. 61112).